OpenStudy (anonymous):

e^xe^(x+1)=1 Part 1: Combine the exponential expression to produce a single exponential expression. Part 2: Convert the equation from part 1 to a logarithmic equation. Part 3: Solve the equation.

OpenStudy (johnweldon1993):

So we have \[\large e^xe^{x + 1} = 1\] right?

OpenStudy (anonymous):

Yea

OpenStudy (johnweldon1993):

We have exponents with the same base (e) so we just add the exponents \[\large e^xe^{x + 1} \Rightarrow e^{x + x + 1} = e^{2x + 1}\] make sense there?

OpenStudy (anonymous):

2x+1=0,x=-1/2

OpenStudy (johnweldon1993):

So what we have is \[\large e^{2x + 1} = 1\]

OpenStudy (anonymous):

\[e ^{0}=1\]

OpenStudy (johnweldon1993):

Now for part 2, we want to change this to a logarithmic equation Well our base being (e) we can take the natural log of both sides (ln) \[\large \ln(e^{2x + 1}) = \ln(1)\] Now remember \(\large \ln(e) = 1\) so we have \[\large 2x + 1 = \ln(1)\]

OpenStudy (anonymous):

Okay I understand that

OpenStudy (johnweldon1993):

Now we just solve \[\large 2x = ln(1) - 1\] We know ln(1) = 0 so \[\large 2x = -1\] And finally after dividing both sides by 2 \[\large x = -1/2\]

OpenStudy (anonymous):

Okay great thanks so much for the help!

OpenStudy (johnweldon1993):

No problem!